### v1

@@ F.4 Analysis of the Second Term in the Generalization Bound
This completes the proof.
@@ F.6 Kernel-Based Hypotheses
For some hypothesis sets, a simpler upper bound can be derived for the ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity of ${\mathscr{H}}$, thereby making Theorems E.5 and F.6 more explicit. We will show this for kernel-based hypotheses. Let $K\colon{\mathscr{X}}\times{\mathscr{X}}\to\mathbb{R}$ be a PDS kernel and let $\Phi\colon{\mathscr{X}}\to\mathbb{H}$ be a feature mapping associated to $K$. We consider kernel-based hypotheses with bounded weight vector: ${\mathscr{H}}_{p}=\left\{(x,y)\mapsto w\cdot\Phi(x,y)\colon w\in\mathbb{R}^{d},\left\|w\right\|_{p}\leq\Lambda_{p}\right\}$, where $\Phi(x,y)=\left(\Phi_{1}(x,y),\ldots,\Phi_{d}(x,y)\right)^{\top}$ is a $d$-dimensional feature vector. A similar analysis can be extended to hypotheses of the form $(x,y)\mapsto w_{y}\cdot\Phi(x,y)$, where $\left\|w_{y}\right\|_{p}\leq\Lambda_{p}$, based on $c$ weight vectors $w_{1},\ldots,w_{c}\in\mathbb{R}^{d}$. The empirical ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity of ${\mathscr{H}}_{p}$ with $p=1$ and $p=2$ can be bounded as follows.
@@
The proofs of Theorems F.7 and F.8 are included in Appendix F.7. Combining Theorem F.7 or Theorem F.8 with Theorem E.5 directly gives the following general margin bounds for kernel-based hypotheses with bounded weighted vectors, respectively.
@@
The first inequality makes use of Hölder’s inequality and the bound on $\left\|w\right\|_{1}$, and the second one follows from the maximal inequality and the fact that a Rademacher variable is 1-sub-Gaussian, and $\sup_{i\in I_{k},y\in{\mathscr{Y}}}\left\|\Phi(x_{i},y)\right\|_{\infty}=r_{k,\infty}$.
See F.8
Proof F.12.
@@
The first inequality makes use of the Cauchy-Schwarz inequality and the bound on $\left\|w\right\|_{2}$, the second follows by Jensen’s inequality, the third by $\operatorname*{\mathbb{E}}[\epsilon_{iy}\epsilon_{jy^{\prime}}]=\operatorname*{\mathbb{E}}[\epsilon_{iy}]\operatorname*{\mathbb{E}}[\epsilon_{jy^{\prime}}]=0$ for $i\neq j$ and $y\neq y^{\prime}$, and the fourth one by $\sup_{i\in I_{k},y\in{\mathscr{Y}}}\left\|\Phi(x_{i},y)\right\|_{2}=r_{k,2}$.

### v2

@@ F.4 Analysis of the Second Term in the Generalization Bound
This completes the proof. ∎
@@ F.6 Kernel-Based Hypotheses
For some hypothesis sets, a simpler upper bound can be derived for the ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity of ${\mathscr{H}}$, thereby making Theorems E.3 and F.2 more explicit. We will show this for kernel-based hypotheses. Let $K\colon{\mathscr{X}}\times{\mathscr{X}}\to\mathbb{R}$ be a PDS kernel and let $\Phi\colon{\mathscr{X}}\to\mathbb{H}$ be a feature mapping associated to $K$. We consider kernel-based hypotheses with bounded weight vector: ${\mathscr{H}}_{p}=\left\{(x,y)\mapsto w\cdot\Phi(x,y)\colon w\in\mathbb{R}^{d},\left\|w\right\|_{p}\leq\Lambda_{p}\right\}$, where $\Phi(x,y)=\left(\Phi_{1}(x,y),\ldots,\Phi_{d}(x,y)\right)^{\top}$ is a $d$-dimensional feature vector. A similar analysis can be extended to hypotheses of the form $(x,y)\mapsto w_{y}\cdot\Phi(x,y)$, where $\left\|w_{y}\right\|_{p}\leq\Lambda_{p}$, based on $c$ weight vectors $w_{1},\ldots,w_{c}\in\mathbb{R}^{d}$. The empirical ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity of ${\mathscr{H}}_{p}$ with $p=1$ and $p=2$ can be bounded as follows.
@@
The proofs of Theorems F.3 and F.4 are included in Appendix F.7. Combining Theorem F.3 or Theorem F.4 with Theorem E.3 directly gives the following general margin bounds for kernel-based hypotheses with bounded weighted vectors, respectively.
@@
The first inequality makes use of Hölder’s inequality and the bound on $\left\|w\right\|_{1}$, and the second one follows from the maximal inequality and the fact that a Rademacher variable is 1-sub-Gaussian, and $\sup_{i\in I_{k},y\in{\mathscr{Y}}}\left\|\Phi(x_{i},y)\right\|_{\infty}=r_{k,\infty}$. ∎
See F.4
Proof.
@@
The first inequality makes use of the Cauchy-Schwarz inequality and the bound on $\left\|w\right\|_{2}$, the second follows by Jensen’s inequality, the third by $\operatorname*{\mathbb{E}}[\epsilon_{iy}\epsilon_{jy^{\prime}}]=\operatorname*{\mathbb{E}}[\epsilon_{iy}]\operatorname*{\mathbb{E}}[\epsilon_{jy^{\prime}}]=0$ for $i\neq j$ and $y\neq y^{\prime}$, and the fourth one by $\sup_{i\in I_{k},y\in{\mathscr{Y}}}\left\|\Phi(x_{i},y)\right\|_{2}=r_{k,2}$. ∎